OpenStudy (wwe123):

if the sum of the first 4 terms of arithmetic progressions p,he sum of the first 8 terms is q and he sum of the first 12 terms r,express 3p+r in terms of q.

hartnn (hartnn):

u applied the formula ? where are u stuck ??

OpenStudy (wwe123):

how express 3p+r in terms of q

hartnn (hartnn):

thats the question.. show your work.

OpenStudy (wwe123):

1 min

hartnn (hartnn):

3p+r = 6*[2a+3d] + 6*[2a+11d] did u get this using formula ?

hartnn (hartnn):

also, 4[2a+7d] = q so, [2a+7d] = q/4

OpenStudy (wwe123):

plz so this ....

hartnn (hartnn):

what ?

OpenStudy (wwe123):

plz solve this ??

hartnn (hartnn):

did u get what i wrote above ?

hartnn (hartnn):

3p+r = 6*[2a+3d] + 6*[2a+11d] = 6*[2a+3d+2a+11d] =6*[4a+14d] = 6*[2(2a+7d)] and 2a+7d was q/4 so, 3p+r = 6*[2(q/4)] = 12q/4 =???

hartnn (hartnn):

got that ^ ?

OpenStudy (wwe123):

answer is 3p+r = 3q

hartnn (hartnn):

yes.

hartnn (hartnn):

12q/4=3q

OpenStudy (wwe123):

thanks :) you are like my Sir

hartnn (hartnn):

welcome ^_^